2. A closed traverse has the folowing data: LINES AB BC 1 BEARING N 30°30E N 75° 30 W S 45°30 W DISTANCES 56.50 m 47.80 m
Answer #1

Lalitude eDepaur-luce +L] N ALLD) I LE +ve D-4ve Lte L - Liase O-W E E+DJ D- lsino [-DJ L=-ve L -ve DE-ve D=tve Lalitude LL]-Le+ve L= +ve D+ve D--ve L= Lalistuole L-ve L-ve De Deparlocc T1 D-ve D= tve (IV. us20 47.8 m s20 gC.7m S6Sm Lm We krouo tocLein 78.727 Lcace O.582 tane p- 4an 12s.27 130. 27 O= 89.57 oplion (c) Posting value L sin 8q2c-78. 727 L= 18.729 L78.73m OpAceo traverse Acct tclorg (142424 Aceo of recct. P0 RC- Ateo o cect= (28.676+78.727) x(60.068+ oc8) ea Trciongk 1+2+3+ ) x 28

LET ME KNOW IF YOU HAVE ANY DOUBT , I WILL BE HAPPY TO HELP.